Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2).

Answer by ewatrrr(24785)

 
Hi,


Line parallel to the graph of 3x - y = -3 OR y = 3x + 3 (green)
Slope of the line parallel is m = 3 (identical slopes)
New line:
y = 3x+ b |using ordered pair (4,-2) to solve for b
-2 = 12 + b
-14 = b
y = 3x -14 (blue)